x+y=7 and 2x+3y=19 can you solve this problem and what should I study to understand this problem.
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

2x+3y+19 has no equal sign.

Answer by Edwin McCravy(20056)



I placed an equal sign where I think you meant to have typed one.

Equation 1:     x + y = 7
Equation 2:   2x + 3y = 19


Solve equation 1 for y:

                x + y = 7
Equation 3:         y = 7 - x

Substitute (7 - x) for y in equation 2:

 
              2x + 3y = 19

        2x + 3(7 - x) = 19

         2x + 21 - 3x = 19

               21 - x = 19

                   -x = -2

                    x = 2

Substitute (2) for x is equation 3:

                    y = 7 - x   
                    
                    y = 7 - (2)

                    y = 5

So the solution is x = 2 and y = 5.

You should go to this website to learn more about this.

http://www.purplemath.com/modules/systlin4.htm

Edwin